Solve

square root of 2y+7 +4=y

Answer by checkley71(8403)

sqrt(2y+7)+4=y
If this is what you mean then:
sqrt(2y+7)=(y-4)
squaring both sides we get:
2y+7=y^2-8y+16
y^2-8y-2y+16-7=0
y^2-10y+9=0
(y-9)(y-1)=0
y-9=0
y=9 answer.
y-1=0
y=1 answer.
proof
sqrt(2*9+7)+4=9
sqrt(18+7+4)=9
sqrt25+4=9
5+4=9